### v1

@@ 7.2 Comparison of the four considered scenarios
In Figure 3, we present the CDF illustrating the downlink rate per user across four examined scenarios utilizing FPA with a parameter of $\alpha=-0.5$, alongside LMMSE beamforming at both BSs and APs, accommodating five users per sector (i.e., $K=180$). The left plot delineates the performance of users within the cell, while the right plot depicts the performance of cell-edge users. Solid lines denote uniform AP placement, while dashed lines denote cell-edge AP placement. Upon scrutiny of the figure, several observations emerge. Firstly, concerning MC-mMIMO deployment, the plots corroborate the significant discrepancy in user performance between those situated within the cell and those positioned at the cell edge. While MC-mMIMO configurations excel in providing remarkable performance to nearby users, they encounter interference limitations when serving users located at the cell edge. Additionally, as anticipated, the FULL-Coop scenario yields the most favorable overall performance. It is noteworthy, however, that there exists an exception where MC-mMIMO outperforms FULL-Coop for cell-interior users in the upper right segment of the CDF. This can be attributed to the detrimental impact of cooperation for users in close proximity to a massive MIMO array, as the additional diversity links are inferior to the direct link with the nearby massive MIMO antenna array. Furthermore, among the schemes exhibiting intermediate performance, HET-noCoop and HORIZONTAL-Coop stand out, with the latter marginally surpassing the former. Regarding AP placement strategies, results indicate that situating APs at the cell edge proves advantageous for cell-edge users but disadvantageous for cell-interior users. This outcome aligns with expectations, as cell-edge AP placement confers benefits to users at the cell edge while somewhat impeding those within the cell. Nevertheless, the degradation experienced by cell-interior users pales in comparison to the gains accrued by cell-edge users, thus reaffirming the intuition that placing APs at the cell edge is preferable. It is also worth noting that, while not explored in this study, employing a more sophisticated association rule between UEs and APs/BSs may mitigate the performance degradation for cell-interior users. This aspect is left for further investigation in future research endeavors.
@@ 7.3 Impact of the beamformers
In Figure 4, our focus is directed towards the two most promising deployments, namely the HORIZONTAL-Coop and the FULL-Coop scenarios, where we present the CDF delineating the downlink rate per user employing FPA with $\alpha=-0.5$. In this analysis, we again assume five users per sector (i.e., $K=180$), and a uniform distribution of APs. The results affirm the well-established hierarchy among the considered beamformers, showcasing that LMMSE achieves superior performance, followed by PZF, and, lastly, MRT combining. Notably, the FULL-Coop deployment surpasses the HORIZONTAL-Coop, with one notable exception observed in the upper portion of the CDF for cell-interior users when MRT is employed. This anomaly can be explained through the limited capacity of MRT beamforming to mitigate interference. In scenarios involving cell-interior users, coercing cooperation with APs and other BSs, as observed in the FULL-Coop scenario, leads to an overall increase in interference. While LMMSE and PZF beamformers well manage this escalation, MRT does not, thereby rendering connecting to other APs/BSs disadvantageous for users benefiting from optimal channel conditions from nearby arrays. This effect could potentially be alleviated by devising more intelligent association rules between UEs and APs/BSs. Moreover, in the lower segment of the CDF, it is discernible that MRT marginally outperforms PZF beamforming, perhaps attributable to the noise amplification effect resulting from projections onto reduced-dimensionality subspaces. Nevertheless, overall, it is evident that the FULL-Coop deployment, employing either LMMSE or PZF beamforming, achieves the most commendable performance across all examined scenarios.
@@ 7.4 Impact of the power allocation strategies
In Figure 5, we examine the influence of the power allocation strategy on the performance of the system. Our analysis is centered on the MC-mMIMO and FULL-Coop scenarios, assuming LMMSE beamforming, a uniform distribution of APs, and 5 users per sector, i.e. we have $K=180$. The figure illustrates the cumulative distribution function (CDF) of the rate per user under the Fixed Power Allocation (FPA) scheme specified in (15), with three different values of $\alpha$: $\alpha=\pm 1$ and $\alpha=0$. Notably, $\alpha=0$ corresponds to uniform power allocation, $\alpha=0.5$ represents partial channel inversion aimed at reducing performance discrepancies among users, while $\alpha=-0.5$ favors network throughput by allocating more power to users with superior channel conditions. Similarly to previous figures, the performance of users within the cell and those at the cell periphery is illustrated separately in two distinct plots. Insightful observations can be made by examining these plots. For the MC-mMIMO scenario, the choice $\alpha=-0.5$ provides the best performance for the cell-inside users (those getting the best channel conditions), and the worse performance for the users at the cell-edge. This experimental result confirms that $\alpha=-0.5$ aims at maximizing the network throughput, neglecting weak users at the expense of network fairness. For the FULL-Coop scenario, instead, a different behavior is obeserved; precisely, the choice $\alpha=-0.5$ provides by far the best performance for cell-inside users, but it also well behaves for the users at the cell-edge. This behavior can be explained by noticing that, thanks to full cooperation, macro-diversity gains provide stable and reasonably good channel conditions to all the uers in the system, and, thus, there is no need to make conservative choices (i.e., taking $\alpha=0.5$) to promote network fairness.
@@
Figure 5: CDF of the rate per user in the MC-mMIMO and FULL-Coop scenarios with three values of $\alpha$ in the FPA strategy, LMMSE beamforming, uniform APs distribution and 5 users per sector, i.e., $K=180$.
@@ 7.5 Impact of the centralized precoding schemes and of the limitation of the fronthaul load
We proceed to examine the influence of the centralized precoding scheme in comparison to the local precoding scheme, concurrently evaluating the impact of enforcing compliance with FH requirements on system performance. In Figures 6 and 7, we present the CDF of the rate per user within the FULL-Coop scenario utilizing FPA with $\alpha=-0.5$, considering cases with 5 and 9 users per sector, respectively. Thus, the total number of users in the system amounts to $K=180$ for Figure 6 and $K=324$ for Figure 7. Our analysis distinguishes between the performance of users located within the cell and those at the cell edge. We employ the PZF beamformer and consider both uniform and cell-edge AP placements. Upon reviewing Figure 6, it becomes evident that the joint PZF beamformer notably outperforms the local PZF beamformer, aligning with our anticipated expectations. Furthermore, adherence to FH requirements introduces a marginal performance degradation in the case of joint PZF beamforming, while the impact on performance in the case of local PZF beamforming is negligible and not discernible in the plots.
Contrasting considerations emerge from Figure 7, which represents the performance of a heavily loaded system. Here, it is observed that meeting the FH constraint results in a noticeable performance decline. However, this degradation is more pronounced in scenarios employing joint centralized beamforming computation compared to those employing local beamforming computation. This discrepancy is rationalized by the fact that beamformers computed at the Central Processing Unit (CPU) necessitate transfer over the FH to the APs/BSs, thereby imposing larger requirements on the FH. Consequently, a greater number of UEs must be disconnected from the APs/BSs to adhere to these requirements, resulting in the observed performance differences. Consequently, the joint PZF beamformer exhibits superior performance over the local PZF beamformer only in unconstrained FH scenarios. Conversely, when considering the limited data rate available on the FH links, it emerges as the least favorable solution. This is a phenomenon that, although counterintuitive, finds justification in the provided explanation. These findings suggest the potential for further research and improvements in the field of data quantization of the fronthaul and about the development of effective association rules between UEs and APs/BSs.
@@ 8 Conclusions
This study has addressed the issue of the simultaneous operation and collaboration between traditional MC-mMIMO and the novel CF-mMIMO network deployments. Given the expected gradual deployment of CF-mMIMO architectures in regions where MC-mMIMO network infrastructures are already in place, the study has examined three potential integration scenarios and compared them against a baseline MC-mMIMO system. The study has presented algorithms for downlink beamforming, AP-UE association, power control, and meeting the FH requirement, conducting a comprehensive performance evaluation through extensive numerical simulations. Various AP deployment strategies have been explored, and performance outcomes for UEs situated within each cell’s core and at the cell periphery have been reported. Overall, the findings indicate that increased levels of integration between APs and BSs are generally advantageous, emphasizing significant performance discrepancies between users at the cell periphery and those in close proximity to BSs. For the former group, collaboration yields substantially higher benefits compared to the latter. The study has also noted that in certain scenarios, the limitation on FH data rates could offset the advantages derived from centralized beamforming, and that in a fully cooperative network environment, employing power control strategies to enhance fairness among users may not always be optimal. In conclusion, it is believed that the issue of collaboration and coexistence between BSs and APs is of significant importance and is likely to attract further research attention in the coming years.
@@
We report here details on the used models for the UE-BS path loss $\text{PL}_{k,\ell}^{(b)}$ and for the UE-AP path loss $\text{PL}_{k,m}^{(a)}$.
With regard to the former term, the path-loss between the $k$-th user and the $\ell$-th BS is expressed as [32, Table 7.4.1-1]
$\text{PL}_{k,\ell}^{(b)}=\max\left(\text{PL}_{k,\ell}^{(b),{\rm LOS}},\widetilde{\text{PL}}_{k,\ell}^{(b)}\right)\,.$ (30)
In (30), the term $\text{PL}_{k,\ell}^{(b),{\rm LOS}}$ is given in Eq. (31), shown on the next page, where $d_{k,\ell,{\rm 3D}}^{(b)}$ and $d_{k,\ell,{\rm 2D}}^{(b)}$ are the 3D and 2D distances between the $k$-th user and the $\ell$-th BS measured in meters, respectively, $f_{c,{\rm GHz}}$ is the carrier frequency measured in GHz and $h_{\rm BS}$ and $h_{{\rm MS},k}$ are the heights of the BSs and user antennas in meters, respectively.
$\text{PL}_{k,\ell}^{(b),{\rm LOS}}[\text{dB}]=\left\{\begin{array}[]{llll}28.0+22\log_{10}\left(d_{k,\ell,{\rm 3D}}^{(b)}\right)+20\log_{10}(f_{c,{\rm GHz}})\quad\quad\text{if}\;10\text{m}\leq d_{k,\ell,{\rm 2D}}^{(b)}\leq\widetilde{d}_{k,\ell,{\rm BP}}^{(b)}\\ 28.0+40\log_{10}\left(d_{k,\ell,{\rm 3D}}^{(b)}\right)+20\log_{10}(f_{c,{\rm GHz}})-9\log_{10}\left[\left(\widetilde{d}_{k,\ell,{\rm BP}}^{(b)}\right)^{2}+\left(h_{\rm BS}-h_{{\rm MS},k}\right)^{2}\right]\;\text{if}\;d_{k,\ell,{\rm 2D}}^{(b)}\geq 5\text{km}\\ \end{array}\right.$ (31)
The breakpoint distance $\widetilde{d}_{k,\ell,{\rm BP}}^{(b)}$ is obtained as
@@
$h_{k,\ell,{\rm E}}^{(b)}=\left\{\begin{array}[]{lll}&1\text{m}\;\;\text{with probability}\;\;p_{k,\ell}^{(b)}\\ &12\text{m}\;\;\text{with probability}\;\;\displaystyle\frac{1-p_{k,\ell}^{(b)}}{3}\\ &15\text{m}\;\;\text{with probability}\;\;\displaystyle\frac{1-p_{k,\ell}^{(b)}}{3}\\ &h_{{\rm MS},k}-1.5\text{m}\;\;\text{with probability}\;\;\displaystyle\frac{1-p_{k,\ell}^{(b)}}{3}\\ \end{array}\right.$ (33)
@@
$C_{k,\ell}^{(b)}=\left\{\begin{array}[]{llll}&\quad\quad\quad 0\quad\quad\quad\quad\quad\quad\quad\text{if}\;h_{{\rm MS},k}<13\text{m}\\ &\left(\displaystyle\frac{h_{{\rm MS},k}-13\text{m}}{10}\right)^{1.5}g_{k,\ell}^{(b)}\;\text{if}\;13\text{m}\leq h_{{\rm MS},k}\leq 23\text{m}\end{array}\right.,$ (34)
$g_{k,\ell}^{(b)}=\left\{\begin{array}[]{llll}&\quad\quad\quad 0\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}\;d_{k,\ell,{\rm 2D}}^{(b)}\leq 18\text{m}\\ &\displaystyle\frac{5}{4}\left(\displaystyle\frac{d_{k,\ell,{\rm 2D}}^{(b)}}{100}\right)^{3}\text{exp}\left(-\frac{d_{k,\ell,{\rm 2D}}^{(b)}}{50}\right)\;\text{if}\;d_{k,\ell,{\rm 2D}}^{(b)}>18\text{m}\end{array}\right.,$ (35)
Finally, in Eq. (30), the term $\widetilde{\text{PL}}_{k,\ell}^{(b)}$ is obtained as
$\begin{array}[]{lll}\widetilde{\text{PL}}_{k,\ell}^{(b)}[\text{dB}]=&13.54+39.08\log_{10}\left(d_{k,\ell,{\rm 3D}}^{(b)}\right)\\ &+20\log_{10}(f_{c,{\rm GHz}})-0.6\left(h_{{\rm MS},k}-1.5\text{m}\right)\,.\end{array}$ (36)
The model in Eq. (30) holds for $1.5\text{m}<h_{{\rm MS},k}<22.5\text{m}$ and $h_{\rm BS}=25$m.
Let us now focus on the path loss model for the UE-AP channel. We use the urban micro cell (UMi) non-line-of-sight (NLOS) model, i.e. the path-loss between the $k$-th user and the $m$-th AP is modeled according to [32, Table 7.4.1-1]:
$\text{PL}_{k,m}^{(a)}=\max\left(\text{PL}_{k,m}^{(a),{\rm LOS}},\widetilde{\text{PL}}_{k,m}^{(a)}\right)\,.$ (37)
In Eq. (37), the term $\text{PL}_{k,m}^{(a),{\rm LOS}}$ is obtained as in Eq. (38), shown on the next page, where $d_{k,m,{\rm 3D}}^{(a)}$ and $d_{k,m,{\rm 2D}}^{(a)}$ are the 3D and 2D distances between the $k$-th user and the $m$-th AP measured in meters, respectively and $h_{\rm AP}$ is the heights of the AP in meters.
@@
$\widetilde{d}_{k,m,{\rm BP}}^{(a)}=4\displaystyle\frac{h_{\rm AP}^{{}^{\prime}}}{h_{{\rm MS},k}^{{}^{\prime}}}\frac{f_{c}}{c},$ (39)
with $h_{\rm AP}^{{}^{\prime}}$ and $h_{{\rm MS},k}^{{}^{\prime}}$ are the effective antenna heights at the AP and the $k$-th user, respectively. The effective antenna heights are computed as $h_{\rm AP}^{{}^{\prime}}=h_{\rm AP}-h_{k,m,{\rm E}}^{(a)}$, $h_{{\rm MS},k}^{{}^{\prime}}=h_{{\rm MS},k}-h_{k,m,{\rm E}}^{(a)}$, where $h_{k,m,{\rm E}}^{(a)}=1$m is the effective environment height. Finally, in Eq. (37), the term $\widetilde{\text{PL}}_{k,m}^{(a)}$ is obtained as
$\begin{array}[]{ll}\widetilde{\text{PL}}_{k,m}^{(a)}[\text{dB}]=&22.4+35.3\log_{10}\left(d_{k,m,{\rm 3D}}^{(a)}\right)\\ &+21.3\log_{10}(f_{c,{\rm GHz}})-0.3\left(h_{{\rm MS},k}-1.5\text{m}\right)\,.\end{array}$ (40)
The model in Eq. (37) holds for $1.5\text{m}<h_{{\rm MS},k}<22.5\text{m}$ and $h_{\rm AP}=10$.

### v2

@@ 7.2 Comparison of the four considered scenarios
Given that the majority of the existing literature on CF-mMIMO systems predominantly addresses the Rayleigh model for small-scale fading, we begin by concentrating on this model, specifically assuming $K_{k,m}^{(a)}=K_{k,\ell}^{(b)}=0$, and defer the discussion of Rician fading to Section 7.7.5. In Figure 4, we present the CDF illustrating the downlink rate per user across four examined scenarios utilizing FPA with a parameter of $\alpha=-0.5$, alongside MMSE beamforming at both BSs and APs, accommodating five users per sector (i.e., $K=180$). The plot on the left delineates the performance of users within the cell, while the plot on the right depicts the performance of cell-edge users. Solid lines denote uniform AP placement, while dashed lines denote cell-edge AP placement. Upon examination of the figure, several observations emerge. Firstly, regarding the deployment of MC-mMIMO, the plots corroborate the significant discrepancy in user performance between those situated within the cell and those positioned at the cell-edge. Although MC-mMIMO configurations excel in providing remarkable performance to nearby users, they encounter interference limitations when serving users located at the cell-edge. Additionally, as anticipated, the FULL-Coop scenario yields the most favorable overall performance. However, it should be noted that there exists an exception in which MC-mMIMO outperforms FULL-Coop for cell-interior users in the upper right segment of the CDF. This can be attributed to the detrimental impact of cooperation for users in close proximity to a massive MIMO array, as the additional diversity links are inferior to the direct link with the nearby massive MIMO antenna array. Furthermore, among the schemes exhibiting intermediate performance, HET-noCoop and HORIZONTAL-Coop stand out, with the latter marginally surpassing the former. Regarding AP placement strategies, the results indicate that placing APs at the cell-edge proves advantageous for cell-edge users but disadvantageous for cell interior users. This outcome aligns with expectations, as cell-edge AP placement confers benefits to users on the cell-edge while somewhat impeding those within the cell. Nevertheless, the degradation experienced by cell-interior users pales in comparison to the gains accrued by cell-edge users, reaffirming the intuition that placing APs at the cell-edge is preferable. It is also worth noting that, although not explored in this study, the use of a more sophisticated association rule between UEs and APs/BSs may mitigate the performance degradation for cell-interior users. This aspect is left for further investigation in future research endeavors.
@@ 7.3 Impact of the beamformers
In Figure 5, our focus is on the two most promising deployments, namely the HORIZONTAL-Coop and FULL-Coop scenarios, where we present the CDF delineating the downlink rate per user employing FPA with $\alpha=-0.5$. In this analysis, again we assume five users per sector (i.e., $K=180$), and a uniform distribution of APs. The results affirm the well-established hierarchy among the considered beamformers, showcasing that MMSE achieves superior performance, followed by PZF, and, lastly, MRT combining. Notably, the FULL-Coop deployment surpasses the HORIZONTAL-Coop, with one notable exception observed in the upper portion of the CDF for cell-interior users when MRT is employed. This anomaly can be explained by the limited capacity of MRT beamforming to mitigate interference. In scenarios involving cell-interior users, coercing cooperation with APs and other BSs, as observed in the FULL-Coop scenario, leads to an overall increase in interference. While MMSE and PZF beamformers manage this escalation well, MRT does not, thereby rendering connecting to other APs/BSs disadvantageous for users benefiting from optimal channel conditions from nearby arrays. This effect could potentially be alleviated by devising more intelligent association rules between UEs and APs/BSs. Moreover, in the lower segment of the CDF, it is discernible that MRT marginally outperforms PZF beamforming, perhaps attributable to the noise amplification effect resulting from projections onto reduced-dimensionality subspaces. However, in general, it is evident that the FULL-Coop deployment, employing either MMSE or PZF beamforming, achieves the most commendable performance across all examined scenarios.
@@ 7.4 Impact of the power allocation strategies
In Figure 6, we examine the influence of the power allocation strategy on the performance of the system. Our analysis is centered on the MC-mMIMO and FULL-Coop scenarios, assuming MMSE beamforming, a uniform distribution of APs, and 5 users per sector, i.e. we have $K=180$. The figure illustrates the cumulative distribution function (CDF) of the rate per user under the Fixed Power Allocation (FPA) scheme specified in (18), with three different values of $\alpha$: $\alpha=\pm 0.5$ and $\alpha=0$. Notably, $\alpha=0$ corresponds to uniform power allocation, $\alpha=0.5$ represents partial channel inversion aimed at reducing performance discrepancies among users, while $\alpha=-0.5$ favors network throughput by allocating more power to users with superior channel conditions. Similarly to previous figures, the performance of users within the cell and those at the cell periphery is illustrated separately in two distinct plots. Insightful observations can be made by examining these plots. For the MC-mMIMO scenario, the choice $\alpha=-0.5$ provides the best performance for the cell-inside users (those getting the best channel conditions) and the worst performance for the users at the cell-edge. This experimental result confirms that $\alpha=-0.5$ aims to maximize network throughput, neglecting weak users at the expense of network fairness. For the FULL-Coop scenario, instead, a different behavior is obeserved; precisely, the choice $\alpha=-0.5$ provides by far the best performance for cell-inside users, but it also behaves well for the users at the cell-edge. This behavior can be explained by noticing that, thanks to full cooperation, macro-diversity gains provide stable and reasonably good channel conditions to all the users in the system, and, thus, there is no need to make conservative choices (i.e., taking $\alpha=0.5$) to promote network fairness.
@@
Figure 6: CDF of the rate per user in the MC-mMIMO and FULL-Coop scenarios with three values of $\alpha$ in the FPA strategy, MMSE beamforming, uniform APs distribution and 5 users per sector, i.e., $K=180$.
Figure 7: CDF of the rate per user in the four considered scenarios with FPA with $\alpha=-0.5$, MMSE beamfoming at the BSs and APs, 5 users per sector, i.e., $K=180$, and uniform APs placement. Continuous lines represent Rayleigh fading (i.e., $K_{k,m}^{(a)}=K_{k,\ell}^{(b)}=0$) and dash-dotted lines represent Rician model detailed in the paper, following the 3GPP 38.901 model.
7.5 Impact of Rician-distributed small-scale fading
In Figure 7, we analyze the performance of the four considered cooperation/co-existence scenarios considering the Rician and Rayleigh channel models described in Section 2.2.2. We assume FPA with a parameter of $\alpha=-0.5$, alongside MMSE beamforming at both BSs and APs, accommodating five users per sector (i.e., $K=180$) and uniform placement of APs. Solid lines denote the performance with the Rayleigh channel model, i.e., $K_{k,m}^{(a)}=K_{k,\ell}^{(b)}=0$, while dash-dotted lines denote the performance with Rician fading, where $K_{k,m}^{(a)}$ and $K_{k,\ell}^{(b)}$ depend on the LOS probability following the model reported in [35]. Examining the figures, it is seen that the relative ranking of the four considered cooperation scenarios does not significantly change; however, the values of the downlink rate when a LOS path is present are larger than those obtained when the LOS path is obstructed. By observing the performance of the MC-mMIMO, we can see that in the lower-left part of the CDF, i.e., the one referring to the worst users, the Rician fading case is outperformed by the Rayleigh one. This behavior can be justified by the higher interference that the worst users experience from the nearby BSs that do not cooperate with the serving one to provide better coverage. In the FULL-Coop scenario, conversely, the Rician fading is always beneficial from the user perspective, but if we focus on the 95%-likely user performance, the two channel models offer very similar performance. Overall, it can be concluded that Rician-distributed channels achieve in most cases better performance than Rayleigh-distributed channels, but the difference is not relevant when focusing on the 95%-likely user performance. Moreover, the relative ranking of the cooperation scenarios considered is mainly preserved. For these reasons, and particularly because rich scattering environments and Rayleigh fading channels serve as the standard benchmark for most simulations in this area of research, the remaining performance results in this study consider the case $K_{k,m}^{(a)}=K_{k,\ell}^{(b)}=0$.
@@ 7.6 Impact of the centralized precoding schemes and of the limitation of the fronthaul load
We proceed to examine the influence of the centralized precoding scheme compared to the local precoding scheme, simultaneously evaluating the impact of enforcing compliance with FH requirements on system performance. In Figures 8 and 9, we present the CDF of the rate per user within the FULL-Coop scenario using FPA with $\alpha=-0.5$, considering cases with 5 and 9 users per sector, respectively. Thus, the total number of users in the system amounts to $K=180$ for Figure 8 and $K=324$ for Figure 9. Our analysis distinguishes between the performance of users located within the cell and those at the cell-edge. We employ the PZF beamformer and consider both uniform and cell-edge AP placements. Upon reviewing Figure 8, it becomes evident that the joint PZF beamformer notably outperforms the local PZF beamformer, aligning with our anticipated expectations. Furthermore, the adherence to the FH requirements introduces a marginal performance degradation in the case of joint PZF beamforming, while the impact on performance in the case of local PZF beamforming is negligible and not discernible in the plots.
Contrasting considerations emerge from Figure 9, which represents the performance of a highly loaded system. Here, it is observed that meeting the FH constraint results in a noticeable performance decline. However, this degradation is more pronounced in scenarios employing joint centralized beamforming computations compared to those employing local beamforming computations. This discrepancy is rationalized by the fact that beamformers computed at the Central Processing Unit (CPU) necessitate transfer over the FH to the APs/BSs, thereby imposing larger requirements on the FH. Consequently, a greater number of UEs must be disconnected from the APs/BSs to comply with these requirements, resulting in the observed performance differences. Consequently, the joint PZF beamformer exhibits superior performance over the local PZF beamformer only in unconstrained FH scenarios. Conversely, when considering the limited data rate available on the FH links, it emerges as the least favorable solution. This is a phenomenon that, although counterintuitive, finds justification in the provided explanation. These findings suggest the potential for further research and improvements in the field of fronthaul data quantization and the development of effective association rules between UEs and APs/BSs.
@@
$\text{PL}_{k,\ell}^{(b),{\rm LOS}}[\text{dB}]=\left\{\begin{array}[]{llll}28.0+22\log_{10}\left(d_{k,\ell,{\rm 3D}}^{(b)}\right)+20\log_{10}(f_{c,{\rm GHz}})\quad\quad\text{if}\;10\;\text{m}\leq d_{k,\ell}^{(b)}\leq\widetilde{d}_{k,\ell,{\rm BP}}^{(b)}\\ 28.0+40\log_{10}\left(d_{k,\ell,{\rm 3D}}^{(b)}\right)+20\log_{10}(f_{c,{\rm GHz}})-9\log_{10}\left[\left(\widetilde{d}_{k,\ell,{\rm BP}}^{(b)}\right)^{2}+\left(h_{\rm BS}-h_{k}\right)^{2}\right]\;\text{if}\;d_{k,\ell}^{(b)}\geq 5\;\text{km}\\ \end{array}\right.$ (38)
@@ 8 Conclusions
This study has addressed the issue of simultaneous operation and collaboration between traditional MC-mMIMO and the novel CF-mMIMO network deployments. Given the expected gradual deployment of CF-mMIMO architectures in regions where MC-mMIMO network infrastructures are already in place, the study has examined three potential integration scenarios and compared them with a baseline MC-mMIMO system. The study has presented algorithms for downlink beamforming, AP-UE association, power control, and meeting the FH requirement, conducting a comprehensive performance evaluation through extensive numerical simulations. Various AP deployment strategies have been explored, and performance outcomes for UEs located within each cell’s core and at the cell periphery have been reported. In general, the findings indicate that increased levels of integration between APs and BSs are generally advantageous, highlighting significant performance discrepancies between users in the cell periphery and those in close proximity to BSs. For the former group, collaboration yields substantially higher benefits compared to the latter. The study has also noted that in certain scenarios, the limitation on FH data rates could offset the advantages derived from centralized beamforming, and that in a fully cooperative network environment, employing power control strategies to enhance fairness among users may not always be optimal. In conclusion, the issue of collaboration and co-existence between BSs and APs is believed to be of significant importance and is likely to attract further research attention in the coming years.
@@
We report here details on the models used for the UE-BS path loss $\text{PL}_{k,\ell}^{(b)}$ and for the UE-AP path loss $\text{PL}_{k,m}^{(a)}$ according to the model in [35, Table 7.4.1-1]. For the two cases, we first evaluate the LOS probability according to Eqs. (33) and (36) and then we select the LOS or NLOS path loss model.
With regard to the term $\text{PL}_{k,\ell}^{(b)}$, in the case of LOS model the path-loss between the $k$-th user and the $\ell$-th BS is expressed in Eq. (38) where $d_{k,\ell,{\rm 3D}}^{(b)}$ and $d_{k,\ell,}^{(b)}$ are the 3D and 2D distances between the $k$-th user and the $\ell$-th BS measured in meters, respectively, $f_{c,{\rm GHz}}$ is the carrier frequency measured in GHz and $h_{\rm BS}$ and $h_{k}$ are the heights of the BSs and user antennas in meters, respectively. The breakpoint distance $\widetilde{d}_{k,\ell,{\rm BP}}^{(b)}$ is obtained as
@@
$h_{k,\ell,{\rm E}}^{(b)}=\left\{\begin{array}[]{lll}1\;\text{m}&\text{with probability}\;\;p_{k,\ell}^{(b)}\\ 12\;\text{m}&\text{with probability}\;\;\displaystyle\frac{1-p_{k,\ell}^{(b)}}{3}\\ 15\;\text{m}&\text{with probability}\;\;\displaystyle\frac{1-p_{k,\ell}^{(b)}}{3}\\ h_{k}-1.5\;\text{m}&\text{with probability}\;\;\displaystyle\frac{1-p_{k,\ell}^{(b)}}{3}\\ \end{array}\right.$ (40)
@@
$C_{k,\ell}^{(b)}=\left\{\begin{array}[]{ll}0&\text{if}\;h_{k}<13\;\text{m}\\ \left(\displaystyle\frac{h_{k}-13\;\text{m}}{10}\right)^{1.5}g_{k,\ell}^{(b)}&\text{if}\;13\;\text{m}\leq h_{k}\leq 23\;\text{m}\end{array}\right.,$ (41)
$g_{k,\ell}^{(b)}=\left\{\begin{array}[]{ll}0&\text{if}\;d_{k,\ell,{\rm 2D}}^{(b)}\leq 18\;\text{m}\\ \displaystyle\frac{5}{4}\left(\displaystyle\frac{d_{k,\ell}^{(b)}}{100}\right)^{3}\text{exp}\left(-\frac{d_{k,\ell}^{(b)}}{50}\right)&\text{if}\;d_{k,\ell}^{(b)}>18\;\text{m}\end{array}\right.,$ (42)
In the case of NLOS model the path-loss between the $k$-th user and the $\ell$-th BS is
$\text{PL}_{k,\ell}^{(b)}=\max\left(\text{PL}_{k,\ell}^{(b),{\rm LOS}},\widetilde{\text{PL}}_{k,\ell}^{(b)}\right)\,,$ (43)
where the term $\widetilde{\text{PL}}_{k,\ell}^{(b)}$ is obtained as
$\begin{array}[]{lll}\widetilde{\text{PL}}_{k,\ell}^{(b)}[\text{dB}]=&13.54+39.08\log_{10}\left(d_{k,\ell,{\rm 3D}}^{(b)}\right)\\ &+20\log_{10}(f_{c,{\rm GHz}})-0.6\left(h_{k}-1.5\;\text{m}\right)\,.\end{array}$ (44)
The model in Eq. (43) holds for $1.5\;\text{m}<h_{k}<22.5\;\text{m}$ and $h_{\rm BS}=25$ m.
Let us now focus on the path loss model for the UE-AP channel. The path-loss between the $k$-th user and the $m$-th AP in the case of LOS is obtained as in Eq. (45), shown on the next page, where $d_{k,m,{\rm 3D}}^{(a)}$ and $d_{k,m}^{(a)}$ are the 3D and 2D distances between the $k$-th user and the $m$-th AP measured in meters, respectively, and $h_{\rm AP}$ is the height of the AP in meters.
@@
$\widetilde{d}_{k,m,{\rm BP}}^{(a)}=4\displaystyle\frac{h_{\rm AP}^{{}^{\prime}}}{h_{k}^{{}^{\prime}}}\frac{f_{c}}{c},$ (46)
with $h_{\rm AP}^{{}^{\prime}}$ and $h_{k}^{{}^{\prime}}$ are the effective antenna heights at the AP and the $k$-th user, respectively. The effective antenna heights are calculated as $h_{\rm AP}^{{}^{\prime}}=h_{\rm AP}-h_{k,m,{\rm E}}^{(a)}$, $h_{k}^{{}^{\prime}}=h_{k}-h_{k,m,{\rm E}}^{(a)}$, where $h_{k,m,{\rm E}}^{(a)}=1$m is the effective environment height. In the case of NLOS, $\text{PL}_{k,m}^{(a)}$ is obtained as
$\text{PL}_{k,m}^{(a)}=\max\left(\text{PL}_{k,m}^{(a),{\rm LOS}},\widetilde{\text{PL}}_{k,m}^{(a)}\right)\,,$ (47)
where the term $\widetilde{\text{PL}}_{k,m}^{(a)}$ is obtained as
$\begin{array}[]{ll}\widetilde{\text{PL}}_{k,m}^{(a)}[\text{dB}]=&22.4+35.3\log_{10}\left(d_{k,m,{\rm 3D}}^{(a)}\right)\\ &+21.3\log_{10}(f_{c,{\rm GHz}})-0.3\left(h_{k}-1.5\;\text{m}\right)\,.\end{array}$ (48)
The model in Eq. (47) holds for $1.5\;\text{m}<h_{k}<22.5\;\text{m}$ and $h_{\rm AP}=10$.